OpenStudy (anonymous):

Solve 2x + 3 = 7 x ≈ –5.19 x ≈ 0.94 x ≈ –0.19 x ≈ 2.81

OpenStudy (anonymous):

LOst -.-

OpenStudy (lgbasallote):

\[2^x + 3 = 7\] or \[2^{x+3} = 7\]

OpenStudy (jiteshmeghwal9):

2x=7-3 2x=4 x=4/2=2

OpenStudy (lgbasallote):

i think it's the second one...

OpenStudy (goformit100):

x ≈ 0.94

OpenStudy (anonymous):

Second one

OpenStudy (lgbasallote):

\[\large 2^{x+3} = 7\] change to log form first. what do you get?

OpenStudy (anonymous):

Thanks but how did you get that?

OpenStudy (lgbasallote):

randomness lol

OpenStudy (anonymous):

Lol

OpenStudy (lgbasallote):

he just picked one and said it...a lot of people do it

OpenStudy (lgbasallote):

anyway...change to log form

OpenStudy (anonymous):

Yeah I know that's why I'm making sure

OpenStudy (anonymous):

I actually did get 0.94 lol could you check it for me

OpenStudy (lgbasallote):

nope. not what im getting

OpenStudy (anonymous):

Whatt :/

OpenStudy (rsadhvika):

you should get negative value for x as 2^x+3 > 7 for any positive value of x

OpenStudy (lgbasallote):

anyway...change to log form...

OpenStudy (anonymous):

Ok so it's x ≈ –0.19

OpenStudy (lgbasallote):

yep

OpenStudy (lgbasallote):

but do you know why?

OpenStudy (anonymous):

Yesss :) I need help with one more than I'll work it out